Write each specification as an absolute value inequality...
6.3 <= h <= 10.3

Answer by stanbon(48545)

Write each specification as an absolute value inequality...
6.3 <= h <= 10.3
----------------------------
Process: (10.3+6.3)/2 = 16.6/2 = 8.3
------------------------------------------
10.3-8.3 = 2
------------------------------
Absolute:
|h-8.3| <= 2
======================
Cheers,
Stan H.